OpenStudy (anonymous):

__ Find the midpoint of PQ if P(7, 0) and Q(8, 0)?

OpenStudy (anonymous):

7.5,0

OpenStudy (anonymous):

is that 15/2, 0?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

for mid point use (x1+x2)/2,(y1+y2)/2

OpenStudy (anonymous):

neg or pos?

OpenStudy (anonymous):

as the result may be

OpenStudy (anonymous):

midpoint of PQ = (7+8/2 , 0+0/2) = (15/2 , 0)

OpenStudy (anonymous):